OpenStudy (anonymous):

Which of the following is the slope between the points (0, -7) and (4, 12)? 19 fourths -19 fourths 4 nineteenths -4 nineteenths

OpenStudy (anonymous):

help plezzzzzzzzzz

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = ( 12-( -7))/(4-0) m = ( 12+7)/(4-0) m = 19/4 So the slope of the line that passes through the two points is 19/4